OpenStudy (anonymous):

how do i find f_xy of f(x,y)=x^3+y^3-xy

OpenStudy (freckles):

first find f_x

OpenStudy (freckles):

To do that pretend y is a constant

OpenStudy (anonymous):

ive found it and f_y. i cant seem to understand the f_xy concept though.

OpenStudy (freckles):

\[(f_x)_y=f_{xy}\]

OpenStudy (freckles):

you find f_x then you find the partial of that w.r.t y

OpenStudy (freckles):

for example \[f(x,y)=e^{3x+2y} \\ f_x=(3+0)e^{3x+2y} =3e^{3x+2y} \\ f_{xy}=(f_x)_y=(3e^{3x+2y})_y=3(0+2)e^{3x+2y}=6e^{3x+2y}\]

OpenStudy (freckles):

so first what did you get for f_x

OpenStudy (anonymous):

i got that f_x = 3x^2-y

OpenStudy (freckles):

\[f=x^3+y^3-xy \\ f_x=3x^2-y \\ f_{xy}=(f_x)_y=(3x^2-y)_y\]

OpenStudy (freckles):

now you find the partial of what you just found w.r.t. to y

OpenStudy (freckles):

\[(3x^2)_y-(y)_y=?\]

OpenStudy (anonymous):

i think its -1

OpenStudy (freckles):

yep

OpenStudy (anonymous):

Awesome! thank you so much!

OpenStudy (freckles):

np